OpenStudy (anonymous):

in meters the sides of a triangle measure 7, 9 and 11. find the perimeter of a similar triangle in which the longest side has a length of 44 meters

OpenStudy (anonymous):

you could set up a propertion for example : 11/9 = 44/x then solve for x, which would be 36 11/7 = 44/x x = 28 then just add up all the sides 44+36+28 so the perimeter is 108

OpenStudy (anonymous):

@bianca101 you need any more help?

OpenStudy (anonymous):

thanks & yes ; in a right triangle the length of a hypotenuse is 20 centimeters and the length of one leg is 16 centimeters find the length of the other leg

OpenStudy (anonymous):

ok so do know the PYTHAGOREAN-THEOREM if not it A^2 + B^2 = C^2

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

sure no problem